OpenStudy (anonymous):

Help with these two math questions? 1.Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. A.–2x – y = –7 B.2x – y = –7 C.2x – y = 11 D. –2x + y = –11 2.Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. A,y – 4 = 2(x + 3) B.y + 4 = 2(x – 3) C.y + 3 = 2(x – 4) D. y – 3 = 2(x + 4)

OpenStudy (anonymous):

use \[y -y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

1. 2x + 7 = y

OpenStudy (anonymous):

which is B is u rearrange it

OpenStudy (anonymous):

2 is the slope, x and y are variables, and 7 is the y intercept.

OpenStudy (anonymous):

2. here are the coordinates that would be if you continued with that slope. -3,4 -2,5 -1,7 0,8 A.y – 4 = 2(x + 3) 5 - 4 = 2[-2-(-3)]